OpenStudy (anonymous):

Solve the system of equations: what is the y-value in the solution? 3x - 5y = -16 and 2x + 5y = 31

OpenStudy (cwrw238):

add the 2 equations to eliminate y then solve resulting equation for x then plug value of x into first equation to find y value

OpenStudy (anonymous):

3x - 5y = -16 2x + 5y = 31 5x + y = 15 Like this?